OpenStudy (anonymous):

The Johnson's are moving to another house. It's takes dad 5 hours to pack all the boxes and it takes 3 hours for mum to pack them. If mom and dad work together, how many hours will it take them to pack all the boxes?

OpenStudy (anonymous):

i have a feeling its \[\frac{ 1 }{ 5 } + \frac{ 1 }{ 3 } = \frac{ 1 }{ x }\]

OpenStudy (anonymous):

my year? If you mean class I'm in Alg 2

OpenStudy (kc_kennylau):

Yep that's correct

OpenStudy (anonymous):

idk what to do now

OpenStudy (anonymous):

@phi could you help me?

OpenStudy (anonymous):

can you add the left side?

OpenStudy (anonymous):

what do you mean? like = 1/8

OpenStudy (kc_kennylau):

No it's not 1/8

OpenStudy (anonymous):

1/3 + 1/5 = 5/15 + 3/15 = 8/15 so 8/15 = 1/x thus, x = 15/8

OpenStudy (kc_kennylau):

\[\frac13=\frac{1\times5}{3\times5}=\frac5{15}\]\[\frac15=\frac{1\times3}{3\times3}=\frac3{15}\]\[\frac13+\frac15=\frac5{15}+\frac3{15}=\frac8{15}\]

OpenStudy (anonymous):

ohhhh u basically just multiplied 1/5 by 3/3 and 1/3 by 5/5 , do i just flip the 8/15 and divide next?

OpenStudy (kc_kennylau):

\[\frac1x=\frac8{15}\]\[\frac1{\frac1x}=\frac1{\frac8{15}}\]\[x=\frac{15}8\]

OpenStudy (kc_kennylau):

You can ignore the second step if you don't understand it.

OpenStudy (anonymous):

ok :), would i just divide 15/8 and reduce the answer so I'd get 1.9 hours?

OpenStudy (kc_kennylau):

Yep

OpenStudy (anonymous):

ok ty

OpenStudy (kc_kennylau):

no problem